# A kinetic model of solar wind acceleration driven by ambipolar electric potential and velocity-space diffusion

### v2

@@ 6 Conclusions and future perspectives
In this paper, we underlined the fundamentals of the existing kinetic exospheric model, discussing possible future improvements of this approach and providing a basis for investigating the influence of electron diffusion on solar wind acceleration. The improvement explored in this study involves redistributing particles within the traditionally considered regions of velocity space — namely ballistic, trapped, and escaping particles — in kinetic exospheric models. This method produces diffused particles capable of reaching perpendicular speeds that would otherwise be inaccessible under strict conservation laws. This constitutes a step toward a model that generates more realistic distribution functions based on the kinetic exospheric theory. In this way, we were able to identify several important features of the kinetic exospheric framework applied to solar wind acceleration:
@@
    A simple implementation of electron diffusion effectively reduces the temperature anisotropy to values comparable to those of the observed fast solar wind at 1 AU.
@@
    This diffusion allows for adjusting the solar wind acceleration without modifying the self-consistent ambipolar electric potential found by the kinetic exospheric model.
@@
    The observed anti-correlation between wind speed and terminal velocity can be reproduced using the kinetic exospheric model with realistic temperatures and exobase levels. This suggests that further analysis of PSP data is necessary to either confirm or rule out the electric potential as the primary acceleration mechanism for the fast solar wind.
@@
    If no indication of the presence of a crossing point between the deduced electric potentials of fast and slow solar winds is found within PSP data, an alternative mechanism — likely velocity-space diffusion of particles via wave-particle interactions — must dominate the fast solar wind acceleration by increasing the flux in the anti-sunward direction.
The extension of the kinetic exospheric model by incorporating an additional acceleration mechanism through diffusion presents a promising approach for improving the model, especially if no evidence of an intersection between the electric potentials of slow and fast solar wind is detected in PSP data. In doing so, it would be possible to determine how much diffusion-dependent net flux is required to explain the observed profiles of the electric potential with their respective terminal velocity. This investigation is left for future work, as the primary objective of this paper is to present the concept of diffusion within the kinetic exospheric model and its ability to provide another source of acceleration (or deceleration).
The primary limitation of the current diffusion implementation is that it relies solely on the initial temperature $T_{0}$ to determine the slope of the diffused population VDF. As a result, the predicted temperatures are overestimated in comparison with observational data. This limitation could be addressed in the future by constraining the slope on the basis of the observed halo or strahl populations and their radial evolution.
Furthermore, greater consistency with the current understanding of the diffusion mechanism could be achieved by explicitly distinguishing between the core and escaping populations, as our current approach considers that the same eVDF defines both core and strahl electrons. This refinement could be implemented following the methodology of Zouganelis et al. 2004, in which a Maxwellian distribution and a Kappa distribution were used to define the core and strahl respectively.

### v3

@@ 6 Conclusions and future perspectives
In this paper we have underlined the fundamentals of the existing kinetic exospheric model, discussing possible future improvements of this approach and providing a basis for investigating the influence of electron diffusion on solar wind acceleration. The improvement explored in this study involves redistributing particles within the traditionally considered regions of velocity space — namely ballistic, trapped, and escaping particles — in kinetic exospheric models. This method produces diffused particles capable of reaching perpendicular speeds that would otherwise be inaccessible under strict conservation laws. This constitutes a step toward a model that generates more realistic distribution functions based on the kinetic exospheric theory. In this way, we were able to identify several important features of the kinetic exospheric framework applied to solar wind acceleration:
@@
    A simple implementation of electron diffusion effectively reduces the temperature anisotropy to values comparable to those of the observed FSW at 1 AU.
@@
    This diffusion enables us to adjust the solar wind acceleration without modifying the self-consistent ambipolar electric potential found by the kinetic exospheric model.
@@
    The observed anticorrelation between wind speed and terminal velocity can be reproduced using the kinetic exospheric model with realistic temperatures and exobase levels. This suggests that further analysis of PSP data is necessary to either confirm or rule out the electric potential as the primary acceleration mechanism for the FSW.
@@
    If no indication of the presence of a crossing point between the deduced electric potentials of the FSW and SSW is found within PSP data, an alternative mechanism — likely a velocity-space diffusion of particles via wave-particle interactions — must dominate the FSW acceleration by increasing the flux in the anti-sunward direction.
The extension of the kinetic exospheric model by incorporating an additional acceleration mechanism through diffusion presents a promising approach for improving the model, especially if no evidence of an intersection between the electric potentials of the SSW and FSW is detected in PSP data. In doing so, it would be possible to determine how much diffusion-dependent net flux is required to explain the observed profiles of the electric potential with their respective terminal velocity. This investigation is left for future work, as the primary objective of this paper is to present the concept of diffusion within the kinetic exospheric model and its ability to provide another source of acceleration (or deceleration).
The primary limitation of the current diffusion implementation is that it relies solely on the initial temperature, $T_{0}$, to determine the slope of the diffused population VDF. As a result, the predicted temperatures are overestimated in comparison with observational data. This limitation could be addressed in the future by constraining the slope on the basis of the observed halo or strahl populations and their radial evolution.
Furthermore, greater consistency with the current understanding of the diffusion mechanism could be achieved by explicitly distinguishing between the core and escaping populations, as our current approach considers that the same electron VDF defines both the core and strahl electrons. This refinement could be implemented following the methodology of Zouganelis et al. 2004, in which a Maxwellian distribution and a Kappa distribution are used to define the core and strahl, respectively.